# Cracking Vector Search Indexes

### v1

@@ 4.3. When to apply build operations?
To mitigate the overhead of constant physical reorganization, which is expensive in high-dimensional IVF indexes, we use a build budget that restricts the fraction of time dedicated to $CRACK$ and $REFINE$ operations. We define $T_{\text{build}}$ as the total measured time spent on previous build operations, which include historical $CRACK$ and $REFINE$. Similarly, $T_{\text{search}}$ represents the cumulative time spent on all past $SEARCH$ operations. When considering a potential build operation, we estimate its execution cost using $\hat{T}_{f}$, which captures the expected time required to perform either a $CRACK$ operation under the current dynamic state ($State_{dyn}$) or a $REFINE$ operation on the local region of the current refine candidate. Here, $f\in\{\textsc{CRACK},\textsc{REFINE}\}$ specifies the build operation considered.
To enforce the budget constraint, we ensure that the total time spent on past and projected build operations remains within a fraction $\alpha$ of the system’s total runtime:
@@
$T_{\text{build}}$ and $T_{\text{search}}$ are directly measured from execution, while $\hat{T}_{f}$ is derived from a predictive cost model, which is explained in the following subsection. The parameter $\alpha$ governs the proportion of total execution time that can be allocated to indexing operations, ensuring a balanced trade-off between indexing efficiency and query performance. This approach allows for a consistent and adaptive mechanism to regulate both cracking and refining decisions under varying workload conditions.
Varying $\alpha$ is a way of expressing an expectation about the number of queries that will arrive. A large $\alpha$ allocates a larger build budget, which is useful if we are optimistic about receiving a large number of queries in the future, thus spending more time to build an index is beneficial. A smaller $\alpha$ restricts the build budget, indicating that we are pessimistic and do not expect many more queries to arrive. To have a balance, in all our experiments, we set $\alpha=0.5$, to ensure a maximum of 50% of the total time at any given point is spent on build operations. This budget ensures that the build operation overhead does not dominate the overall execution time. As more queries arrive, CrackIVF will progressively have a larger budget to apply for physical reorganization operations. In this way, we can remain near the optimal minimum of cumulative time across all query scales. Finally, there is a minimum size constraint on the number of buffered cracks before a $CRACK$ is executed. Specifically, in our implementation, we set that minimum to 20% of the current index size, $|C_{buffered}|>C_{min}$, where $C_{min}=0.2*|C|$ (Line 19 Algorithm 1). This helps avoid situations where a $REFINE$ operation applied on the current index achieves similar performance improvements, while allowing us to keeping the current index size without increasing it.
Predictive Cost Model
The cost estimation of both CRACK and REFINE operations happens after every incoming query. Thus, it is crucial to maintain low latency when making a prediction while at the same time maintaining an accurate enough estimate to satisfy our usage needs. To achieve this balance, we employ a simple first-order linear model and fit it using multivariate regression (Jain, 1990).
Each build function $f$ comprises a sequence of procedures, or kernels, indexed by $i$ and executed sequentially, as shown in Algorithm 2 and Algorithm 3. Among these, updating the index state is negligible, and the following kernels are the dominant factors of each function’s execution time:
@@
Since the above kernels in our implementation execute sequentially without overlap, the total execution time $T_{f}$ is simply the sum of the execution times of its kernels.
@@
Each kernel $i$ contributes a latency $T_{f,i}$, which we model as a linear function of its two dominant cost factors: computational complexity and data movement. Specifically, for each kernel $i$ in function $f$, we train the following simple predictive model:
@@
where $C_{f,i}$ represents the dominant computational complexity of kernel $i$, and $D_{f,i}$ denotes the dominant data movement cost. The terms $w_{f,i}^{(1)}$ and $w_{f,i}^{(2)}$ are learned coefficients corresponding to computation and data movement, respectively. Finally, $b_{f,i}$ is a learned parameter that accounts for fixed kernel execution overheads.
We gather measurement with micro-benchmarks for each kernel, by varying input parameters and measurig execution latency to fit a regression model, learning $w_{f,i}^{(1)},w_{f,i}^{(2)},b_{f,i}$.
Our approach eliminates the need for handcrafted analytical performance models that map the kernels into the number of cycles per operation and are tied to specific implementations and hardware configurations, making them hard to adapt or extend. With our method, switching hardware or optimizing kernels only requires re-running microbenchmarks and refitting the linear model. Another advantage is the low computational cost of inference, enabling continuous use during query execution. However, modeling execution time as a simple linear function has limitations. While it captures dominant compute and memory trends, it overlooks system-level nonlinearities. Factors like caching effects, NUMA constraints, SIMD optimizations (e.g., in FAISS), and thread contention introduce discrete performance variations that a linear model cannot fully express.
@@
Get Local Region* $|C_{\text{buffered}}|\times|C|$ $|P_{\text{local}}|\times d$
@@
Table 2. Features we used for each kernel. * For REFINE, $C_{\text{buffered}}=1$
Nevertheless, our empirical results show that the cost model performs well across all benchmarks. Importantly, we always use the true measured times $T_{\text{build}}$ and $T_{\text{search}}$ in our budget constraint Equation 8. This ensuring any overestimation by the predictive model is corrected after execution, and an underestimation will hold any future build operations until sufficient queries arrive, increasing the total budget. In our testing, the model achieved moderate to high $R^{2}$, explaining between 40 and 95% of the total variation, with the $RMSE$ remaining in the millisecond to low second range. We find this to be acceptable performance from our model since we benefit from low inference latency, and the cost budget-based control mechanism works effectively on all datasets, as we show in our experiments.
@@ 5. Experiments
5.1. Experimental Setup
We run on a dual-socket AMD EPYC 7V13 system (128 cores, 3.7 GHz, 512GB RAM) from the AMD-ETHZ HACC cluster (Moya et al., 2023a, b, 2024). Unless otherwise stated, experiments use 16 cores, the number at which the memory bandwidth saturates, and use an equal query batch size to benefit from inter-query parallelism.
Table 3. AV-Tree vs. CrackIVF
Cumulative Time at i-th Query (seconds)
$10^{0}$ $10^{1}$ $10^{2}$ $10^{3}$ $10^{4}$ $10^{5}$ (Total)
AV-Tree 0.078 0.747 7.601 75.369 825.709 1264.175
CrackIVF 1.107 1.166 1.738 15.983 98.295 516.881
Datasets: We evaluate on standard ANN search benchmarks for text and image-based retrieval, GloVe(Pennington et al., 2014), SIFT(Jégou et al., 2011), DEEP(Babenko and Lempitsky, 2016), and Last.fm(Bertin-Mahieux et al., 2011). The latter, being an example with a high skew, is used for music recommendation. The SIFT1M and SIFT10M, contain 128-dimensional SIFT descriptors, while DEEP10M consists of 96-dimensional embeddings from a fully connected GoogLeNet model (Szegedy et al., 2015). These datasets are subsets of the original billion-scale collections, with 1M and 10M referring to the number of data points in each slice. Both use L2 distance as a metric. GloVe, sourced from ANN-Benchmarks (Aumüller et al., 2020), represents text retrieval. It contains 1.18M varying dimensional data points and uses cosine similarity, which we implement by l2-normalizing and computing inner product (IP) similarity. Last.fm is also taken from (Aumüller et al., 2020); it contains 64-dimensional vectors and uses the IP similarity metric. We found that the provided query set is highly skewed, allowing us to test how CrackIVF behaves in such a scenario. Last.fm provides 50k unique queries, while the other datasets provide 10K unique queries. Thus, we replicate them until we match our target scales (up to 1M).
Baselines: Our comparisons cover both exact nearest neighbor (ENN) and approximate nearest neighbor (ANN) search approaches. For ENN, we use Brute Force, i.e., a linear scan of the data and AV-Tree (Lampropoulos et al., 2023), as a cracking baseline. For ANN search, we compare against FAISS IVF indexes, which in our figures are indicated as IVFFlat, with the number of partitions chosen in parentheses, e.g., IVFFlat (1000). We vary the number of partitions from 1,000 to 16,000, following FAISS guidelines (Facebook Research, 2025), which recommend this range for our dataset scales. Our method, CrackIVF is initialized with 100 partitions to maintain near brute force startup time.
All experiments have the same fixed parameters, with values set as explained in (Sections 4.2, 4.3). The $\alpha$ parameter to control the budget remains fixed to $0.5$. Allowing, a build budget equal to 50% of the total time. Finally, for the partition-based indices, we have CrackIVF and IVFFlat baselines, where the $nprobe$ is set to always achieve 90–95% Recall@10. The QPS-Recall plots show the Pareto frontier when varying $nprobe$ across all indexes.
@@ 5.2. Comparison with AV-Tree:
To compare with AV-Tree (Lampropoulos et al., 2023) fairly, we set our CrackIVF to run with a single thread and make sure we probe enough partitions to get an average of more than $99\%$ recall across all query scales. We set the cracking threshold of AV-Tree to 128, which we find gives the best total runtime. We show the cumulative time for a trace of $100$K queries over the SIFT 1M dataset in Table 3. Unlike ours, AV-Tree does not pre-build the index, therefore, it has an advantage over us for the first 10 queries. However, our method processes each query faster and outperforms AV-Tree by a large margin after the first few queries. The cumulative time of AV-Tree increases at a slower rate with more queries, meaning that the query time is getting faster, and AV-Tree is improving over time. Nevertheless, our CrackIVF is 2.45x faster than AV-Tree in the end, even on a single-threaded ENN search. This shows the effectiveness of ANN approaches, where a less than 0.01% Recall error leads to a large gain in performance. For the rest of the experiments, we focus on the ANN search setting.
@@ 5.3. Does CrackIVF improve over time?
To measure how CrackIVF improves over time, we provide two graphs for each of the datasets and group them in Figure 3. The QPS vs. Recall plots measure the maximum achievable Queries Per Second (QPS) across varying Recall targets. This is the trade-off under which the ANN search operates. Having a large gain in search speed (QPS) for a small decrease in search quality (Recall). By the end of the 1 million query trace, indicated by CrackIVF-END (orange line), we are consistently at the Pareto frontier of QPS-Recall across on all datasets. On SIFT10M and Last.fm, CrackIVF is by far the best-performing index. Specifically on Last.fm, CrackIVF with 1255 final partitions, achieves 50% higher QPS for Recall@10 = 0.9, than a 10x larger IVFFlat(16000) index. We attribute this to CrackIVF’s approach of only allocating new cracks based on the query distribution, and since Last.fm is the dataset with the highest skew, where only  30% of the vector spaces is accessed (Figure 2(b)), we can fully partition that  30% of the space with a relatively small number of cracks to achieve the best performance.
In Figure 3, we also plot the Response Time per query batch across the entire 1 million query trace. We can see that CrackIVF consistently improves over time, showing drops in response time after incremental reorganization operations. The outliers can all be attributed to these $CRACK$ and $REFINE$ reorganization operations. This is evidence for having control mechanisms to ensure expensive operations are only applied when necessary and why CrackIVF relies on buffering crack operations. The downstream of buffering and committing a large number of multiple cracks at once is only visible in one dataset, Last.dm, where it creates a step-like response in the data, as seen in Figure 3(f). This sudden improvement in performance when adding a lot of cracks can also be seen in Figure 2(a), when going from 1000 to 5000 partitions.
@@ 5.4. Does CrackIVF Minimize Cumulative Time?
Our core motivation in this work has been to provide an index that can be deployed across data lakes under scenarios where the number of queries that will arrive for each dataset and utilize the index is unknown. The ideal index for such a scenario should be able to minimize the cumulative time spent on both build and search operations in cases where a dataset may receive a few queries as well as millions.
@@
We evaluate this specific aspect of index performance, across all datasets, with the cumulative time experiments, shown in Figure 4. For our measurements, we run a trace of 1 million queries for each dataset. In all baseline methods, we include the upfront build cost as the time before they answer the first query. The smaller IVFFlat (1000) indexes have relatively low startup costs but do not scale well past  10k queries since their cumulative time rises sharply due to having higher query batch response times than the alternative. The larger indexes, beyond 5000 partitions, have a huge upfront build cost, which may pay off but only if a dataset receives ¿1 million queries. The “no index“ approach of the Brute-force method has the minimum initialization time and is thus the best choice when the number of queries is low (¡100).
CrackIVF consistently achieves near-minimum cumulative time across all query scales. It achieves up to 1000x lower initialization time than larger indexes (Figure 4(d)) and can consistently answer 100K to 1M queries before comparable indexes finish building. For example on SIFT10M (Figure 4(b)), the initialization time cost is 10x lower than IndexFlat(5000) and 100x lower than IndexFlat(16000), yet it processes 100K and 1M queries respectively, before they baselines indexes complete construction, while at the same time converging to higher QPS-Recall performance. Similar results can be observed in all other datasets.
The fact that CrackIVF achieves the Pareto optimal result in QPS-Recall by the end of our trace means that, by extrapolation, it will always maintain the minimum cumulative time across all query scales beyond 1M queries.

### v2

@@ 4.3. When to apply build operations?
To mitigate the overhead of constant physical reorganization, we use a build budget to restrict the fraction of time dedicated to $CRACK$ and $REFINE$ operations. We define $T_{\text{build}}$ as the total measured time spent on previous build operations. Similarly, $T_{\text{search}}$ represents the cumulative time spent on all past search operations. When considering a potential build operation, we estimate its execution cost using $\hat{T}_{f}$, which captures the expected time required to perform a $CRACK$ operation in the current dynamic state or a $REFINE$ operation in the true state. The build operation considered is specified by $f\in\{\textsc{CRACK},\textsc{REFINE}\}$. To enforce the budget constraint, the total time spent on past build operations in addition to the estimated time if the current one is executed must remain within a fraction $\alpha$ of the system’s total runtime:
@@
$T_{\text{build}}$ and $T_{\text{search}}$ are directly measured from the execution time, while $\hat{T}_{f}$ is estimated using a predictive cost model. The parameter $\alpha$ defines the proportion of total execution time that can be allocated to indexing operations. We set $\alpha=0.5$, which ensures a maximum of 50% of the total time at any given point is spent on build operations. As more queries arrive, CrackIVF will progressively have a larger budget to apply for physical reorganization operations. In this way, we can remain near the optimal minimum of cumulative time across all query scales. Finally, there is a minimum size constraint on the number of buffered cracks before a $CRACK$ is executed. The idea comes from Figure 2(a), where if a REFINE can achieve similar performance improvement at the current index size, it is better to wait until more cracks have been buffered, so that the performance improvement from index growth is substantial. Specifically, in our implementation, we set that minimum to 20% of the current index size, $C_{min}=0.2\times|C|$ (Line 19 Algorithm 1).
Predictive Cost Model: Estimating the execution time of the CRACK and REFINE operations happens after every incoming query. Thus, we must maintain low latency when making a prediction, while have accurate enough cost estimates to correctly use the budget constraint Eq. 6. For this, we use a simple first-order linear model fitted using multivariate regression (Jain, 1991). Each build function $f$ comprises a sequence of procedures, or kernels, indexed by $i$ and executed sequentially, as shown in Algorithm 2 and Algorithm 3. The execution time is dominated by the following kernels:
@@
These kernels are executed sequentially so the total execution time $T_{f}$ is simply the sum of kernel execution times. Each kernel $i$ contributes a latency $T_{f,i}$, which we model as a linear function of its two dominant cost factors: computational complexity and data movement. For each, we learn the following predictive model:
@@
where $C_{f,i}$ represents the dominant computational complexity of kernel $i$, and $D_{f,i}$ denotes the dominant data movement cost. The terms $w_{f,i}^{(1)}$ and $w_{f,i}^{(2)}$ are learned coefficients for computation and data movement and $b_{f,i}$ is a learned parameter that accounts for fixed overheads. We gather measurement with micro-benchmarks for each kernel, by varying input parameters and measurig execution latency to fit a regression model, learning $w_{f,i}^{(1)},w_{f,i}^{(2)},b_{f,i}$.
Table 2. Features used for each kernel. For REFINE, $C_{\text{buffered}}=1$
@@
Get Local Region $|C_{\text{buffered}}|\times|C|$ $|P_{\text{local}}|\times d$
@@
Our approach removes the need for exact analytical models coupled to specific implementations and hardware. Switching hardware requires re-running micro-benchmarks and refitting the linear model. The main limitation is that execution time is modeled as a linear function. While this captures dominant compute and memory trends, it overlooks system-level nonlinearities such as caching effects, NUMA constraints, and SIMD optimizations. Despite this, our empirical results show that the cost model performs well across all benchmarks. Importantly, we always use the true measured times $T_{\text{build}}$ and $T_{\text{search}}$ in the budget constraint (Equation 6), which help correct over and under estimations. In our tests, the model achieved moderate to high $R^{2}$ scores, explaining 40–95% of the total variance, with RMSE ranging from milliseconds to a few seconds.
@@ 5. Experiments
We run on a dual-socket AMD EPYC 7V13 system (128 cores, 3.7 GHz, 512GB RAM). All experiments, except for Table 3, use 16 cores and a batch size of 16, the point at which memory bandwidth saturates. IVF and CrackIVF share the same SEARCH operation in FAISS and benefit equally from inter-query parallelism.
We evaluate on standard ANN search benchmarks, GloVe(Pennington et al., 2014), SIFT(Jégou et al., 2011), DEEP(Babenko and Lempitsky, 2016), and Last.fm(Bertin-Mahieux et al., 2011). SIFT1M and SIFT10M contain 128-dimensional SIFT descriptors; DEEP10M consists of 96-dimensional embeddings, with 1M and 10M being the number of data points in each slice. Both use L2 distance as a metric. GloVe contains 1.18M varying dimensional embeddings and uses cosine similarity, which we implement by l2-normalizing and computing the inner product (IP). Last.fm, contains 64-dimensional vectors and IP metric. The Last.fm query set is highly skewed, allowing us to test CrackIVF in such cases. Last.fm provides 50k unique queries, while the other datasets provide 10K unique queries. We replicate them until we match our target scales (up to 1M).
For ENN, we compare against Brute Force, and AV-Tree (Lampropoulos et al., 2023), which is a cracking baseline. For ANN search, we compare against FAISS IVFFlat indexes with the number of partitions chosen shown in parentheses, e.g., IVFFlat (1000). We vary the number of partitions following FAISS guidelines (Facebook Research, 2025), for our dataset scales. For CrackIVF, we use fixed parameters values for the control mechanisms, set as explained in (Sections 4.2, 4.3). We consider the index converged when the crack decision rules stop triggering and refines keep targeting the same region without effect, thus the index can no longer grow in size, nor improve at the current size. For CrackIVF and IVFFlat, nprobenprobe is set to always achieve 90–95% Recall@10 in ANN. CrackIVF is initialized with 100 partitions on all experiments. We set this parameter by choosing the largest possible starting partitions that still give minimal startup cost. Because of FAISS parallelism, at 16 threads, the startup time for 100 partitions is identical or faster than starting with 1 partition. On a single thread for Table 3, the difference between them is only $\approx$120 ms.
@@ 5.1. Comparison with AV-Tree:
Table 3. AV-Tree vs. CrackIVF. Cracking-based baselines.
Cumulative Time at i-th Query (seconds)
$10^{0}$ $10^{1}$ $10^{2}$ $10^{3}$ $10^{4}$ $10^{5}$ $10^{6}$
IVF-1K 12.57 12.59 12.76 14.48 32.10 205.40 1938.60
IVF-5K 178.17 178.18 178.35 179.93 195.94 354.63 1943.02
AV-Tree 0.08 0.75 7.48 75.43 830.0 1274.85 5721.29
CrackIVF 1.10 1.16 1.73 13.18 55.33 233.11 1531.40
To the best of our knowledge, AV-tree (Lampropoulos et al., 2023) is the only similarity search index, whose underlying idea is based on cracking. It is a tree-based ENN cracking index targeting short-lived data up to one thousand queries. Unlike AV-Tree, CrackIVF is an IVF-based ANN cracking index, implemented around FAISS, offering parallelism, and targeting high dimensional RAG applications which could reach much larger query volumes. As AV-Tree is single-threaded and an ENN index, to compare fairly, we set CrackIVF and the IVF baselines to run with a single thread and make sure we probe enough partitions to get an average of more than $99\%$ recall across all query scales. We set the cracking threshold of AV-Tree to 128, which we find gives the best total runtime. We show the cumulative time for a trace of 11M queries over the SIFT 1M dataset in Table 3. At the start, the $10^{0}$-th query time includes the index construction costs of CrackIVF and IVF baselines. CrackIVF is initialized as a small 100 partition index which grows as more queries arrive, whereas AV-Tree does not pre-build any index. In this experiment, AV-Tree has an advantage over CrackIVF for the first 10 queries, as it can start answering immediately. However, CrackIVF processes each query faster and outperforms AV-Tree by a large margin after the first few queries. The cumulative time of AV-Tree increases at a slower rate with more queries, showing AV-Tree is improving over time. Our CrackIVF is 3.7x faster than AV-Tree in the end of the trace. As cracking baselines, both CrackIVF and AV-Tree, vastly outperform IVF indexes on the low query scale. AV-Tree continues to outperform up to 1000 queries, matching original claims (Lampropoulos et al., 2023). CrackIVF performance benefits continue even at 1 million queries. In the following experiments we focus on ANN.
@@ 5.2. Does CrackIVF improve over time?
We measure how CrackIVF improves over time in Figure 3. QPS vs. Recall plots show the search performance trade-off in which ANN search operates under. In our experiments, by the end of the query trace, CrackIVF-END is consistently at or near the Pareto frontier across all tested datasets. On SIFT10M and Last.fm, CrackIVF is by far the best-performing index. Specifically on Last.fm, CrackIVF with 1264 final partitions, achieves $\approx$50% higher QPS for Recall@10 = 0.9, than any other index. We attribute this to CrackIVF’s approach of allocating new cracks following the query distribution. Since Last.fm is the dataset with the highest skew (Figure 2(b)), we can fully partition the accessed space with a small number of cracks, achieving the best performance. In Figure 3, we also plot the response time per query batch across the entire query trace. CrackIVF search performance improves over time after incremental reorganization operations. The vast majority of outliers are due to the individual $CRACK$ and $REFINE$ reorganization operations. In our experiments, $CRACK$ operations are less than 0.03% and $REFINE$ less than 0.6% of the query trace.
@@ 5.3. Does CrackIVF Minimize Cumulative Time?
We run a trace of 1 million queries for each dataset and include the upfront build cost for all baselines as the time before they answer the first query (Figure 4). The smaller 1000 partition IVF indexes have relatively low startup costs but do not scale well past $\approx$10k queries. Larger indexes have a huge upfront build cost, which only pays off if a dataset receives ¿1 million queries. Brute-force has the minimum initialization time, making it the better choice when the number of queries is low (¡100). In our experiments, CrackIVF achieves near-minimum cumulative time across all query scales. It achieves almost 1000x lower initialization time than larger indexes (Figure 4(d)) and can answer 100K to 1M queries before comparable indexes finish building. For example on SIFT10M (Figure 4(b)), the initialization time cost is $\approx$100x lower than IndexFlat(5000) and almost 1000x lower than IndexFlat(16000), yet CrackIVF answers 100k and 1M queries respectively, before the indexes finish building.
@@
5.4. Does CrackIVF Reduce Total Distance Computations for Build Operations?
IVF indexes incur full indexing cost upfront over the entire vector space. Indexing large datasets (100M–1B+) in reasonable time requires GPUs, but even then, K-means training can take hours due to the increasing number of training points and partitions, which have a multiplicative effect in the number of pairwise distance computations (see Table 2 K-means Compute). CrackIVF amortizes this cost over time using $REFINE$ operations, which apply K-means locally on small subsets of points and partitions. As a result, total distance computations are spread across time and space and can potentially be handled on cheaper hardware (e.g., CPUs).
Figure 5 shows the cumulative distance computations during index build, excluding search operations. CrackIVF pays this cost gradually, and after the index converges, the final distance computations can be orders of magnitude lower than larger IVF indexes (Figure 5(a)5(f)). Despite having comparable or better performance (Figure 3(a)3(k)). Convergence occurs when the arrival of new crack candidates stops and refines do not trigger on new regions, example is shown in ablation. On Glove, distance computations increase faster due to $REFINE$ regions containing more points and more cracks. The median $REFINE$ in Glove100 covers 4.784.78% of the dataset with 512 cracks, while the median $REFINE$ in SIFT1M is only $0.63\%$ of the total points across 64 cracks. Nonetheless, the total cost remains amortized over time and is still almost an order of magnitude lower than that of the largest IVF index, while achieving similar final QPS-Recall performance. Finally, CRACK is dominated by data movement, while REFINE is mostly compute (Table 2). Even though CRACK is infrequent ($0.03\%$ of trace Figure 3), its overhead is not fully captured by distance computations alone. This is why we use fixed hardware and cumulative time plots, to compare the total time across all operations, including search, in Figure 4.
(a) SIFT1M
(b) SIFT10M
(c) Last.fm
(d) Glove50
(e) Glove100
(f) Deep10M
Figure 5. Cumulative distance computations for index build.
(a) WHERE=OFF
(b) WHERE=ON
Figure 6. Effect of removing the “where” control mechanism
(a) Cumulative Time (OFF)
(b) Time Breakdown (OFF)
(c) Cumulative Time (ON)
(d) Time Breakdown (ON)
Figure 7. Effect of removing the “when” control mechanism